OpenStudy (anonymous):

6th Grade Math I will Fan and Medal the person who helps.

RhondaSommer (rhondasommer):

Im here :D whats your question?

OpenStudy (anonymous):

I'll put the question up in a second. :)

RhondaSommer (rhondasommer):

:D okay :D

OpenStudy (anonymous):

I think it is C.

OpenStudy (anonymous):

But I have to pick two answers.

OpenStudy (er.mohd.amir):

just try cross multiply get II ans.

OpenStudy (baru):

you picked \[r=\frac{14}{4}\] can you do a simple algebraic re-arrangement by multiplying both sides by 4?

OpenStudy (baru):

both sides of the equation i mean

OpenStudy (anonymous):

What do you mean, sorry i'm really bad at math xd

OpenStudy (baru):

we have \[r=\frac{14}{4}\] multiply the left and the right side of the equation by 4. what do you get?

OpenStudy (anonymous):

56 and 16? o-o

OpenStudy (anonymous):

Because I multiplied 14 and 4..? Is that what I'm supposed to do? xD

OpenStudy (baru):

multiply both sides of \[r=\frac{14}{4}\] by 4 we get \[4 \times r=\frac{14 \times 4}{4}\]

OpenStudy (baru):

can you simplify that

OpenStudy (anonymous):

Ohhhhhhhhhhhh

OpenStudy (anonymous):

probably

OpenStudy (baru):

give it a try, what are you left with?

OpenStudy (anonymous):

I am left with = 7/2 = 3 1/2

OpenStudy (baru):

can you do this \[\frac{4}{4} = ?\]

OpenStudy (anonymous):

yeah, 1

OpenStudy (baru):

lets look at the right side \[ \frac{14 \times 4}{4}\] it is the same as \[ 14 \times \frac{ 4}{4}\] so \[ 14 \times \frac{ 4}{4}=?\]

OpenStudy (anonymous):

It is 14.

OpenStudy (anonymous):

14*4/4=14?

OpenStudy (baru):

yes!

OpenStudy (anonymous):

So is that my answer? xD

OpenStudy (baru):

so the right side of \[4 \times r=\frac{14 \times 4}{4}\] becomes just 14 leaving us with \[4 \times r=14\]

OpenStudy (baru):

does that look like any of your options?

OpenStudy (anonymous):

Um, no can you check my options?

OpenStudy (baru):

well, i'll just tell you that it does look like one of your options. you will have to look it up :)

OpenStudy (anonymous):

I'll just go with A and D, thanks for the help.